Find v(w + 3) for v(w) = 2w - 1
Answer by stanbon(75887)

Find v(w + 3) for v(w) = 2w - 1
---------------------------
v(w+3) = 2(w+3)-1 = 2w+6-1 = 2w + 5
------------------------
Cheers,
Stan H.
------------